Two blocks are connected by a string that is wrapped around a pulley wheel. The light block with mass 5.3 kg sits a flat table; the heavier block, which is twice as massive, is suspended below the pulley. There is friction acting between the lighter block and the table with a coefficient of friction 0.79. The system is released from rest. What is the speed of the blocks after the bottom block has fallen 4.2 meters in units of meters/second? You can ignore the mass of the string and the pulley wheel.
Answer #1

o a T my % AT a Along vertical! al mig mi=5.3kg M2=2m. m₂ =2x5.3 kg M2 =10.6 kg m la Tis tension in the string. a is accelera41:0326 = 5:30 a For Mr. AT Im La Mag a = M g-T= mza > (10.6x9.8) – T=10.69 103-88-T= 10.69 from equation 6 and ② 7-41.0326 +